OpenStudy (yacoub1993):

Factor the following expression. Show you work with steps. --> 198q3r2 – 184q2r2 + 18qr2

OpenStudy (yacoub1993):

@DSS @dss08a @dpasingh @caitlinnr14 @mathessentials @Hero @helperoc101 @hba @Aiko @Ahmad1 Help me Please....

OpenStudy (yacoub1993):

\[198q ^{3} r ^{2}–184q ^{2} r ^{2}+18qr ^{2}\]

OpenStudy (anonymous):

\[198q^3r^2 – 184q^2r^2 + 18qr^2 = qr^2 (198q^2-184q+18)\] \[= 2qr^2 (99q^2-92q+9)\] \[= 2qr^2 (99q^2-92q+9)= 2qr^2 (99q^2-11q-81q+9)\] \[= 2qr^2[11q (9q-1)-9(9q-1)]= 2qr^2(11q-9) (9q-1)\] Thus \[2qr^2(11q-9) (9q-1) \] are the required factors of the given expression.

OpenStudy (anonymous):

@Yacoub1993

OpenStudy (anonymous):

@Yacoub1993 Your question has been solved. Pls visit the page for solution.

OpenStudy (yacoub1993):

woow i wasnt expecting that anyways thnxs @dpasingh